### v1

# Searching for Periodicity in FRB 20240114A

Abstract
FRB 20240114A is extraordinarily active, and therefore presents an opportunity to search for the periodicity predicted by magnetar models of Fast Radio Bursts (FRB). Zhang et al. (2025) observed 11,553 bursts, including 3196 on MJD 60381 (March 12, 2024). We find no significant peak in the periodogram of those bursts, which occur within 15628 s, short enough that even with a characteristic slowing age of 1 year a signal with period $\geq 0.1\,$s it would not significantly dephase within the observation. Introducing modulation artificially shows that an amplitude of 0.15 would have been detected robustly. Implications for magnetar models of FRB are briefly considered.
radio transient sources
I Introduction
The most popular model of Fast Radio Bursts (FRB) (Zhang, 2023) attributes them to magnetars, hypermagnetized neutron stars. Magnetars were first proposed (Katz, 1982; Duncan & Thompson, 1992) as models of Soft Gamma Repeaters (SGR; initially considered to be gamma-ray bursts), a model that was verified when rapid spindown between outbursts confirmed their extraordinary magnetic fields. This model makes one obvious prediction: that FRB, or the level of FRB activity are periodic with the neutron star’s rotational period. This prediction has never been verified for any FRB, despite evidence Niu, Li, Yang et al. (2025); Wang, Lan, Zhao et al. (2025) that at least one FRB is inside an expanding supernova remnant, as would be exptected in a magnetar model.
The extreme activity and many detected bursts of FRB 20240114A, including 3196 within 15628 s on the single day March 12, 2024 (Zhang et al., 2025), present a favorable opportunity for a search. We report the failure of that search.
II Types of periodicity
We distinguish two types of periodicity:
II.1 RRAT periodicity
RRAT (Rotating RAdio Transients) emit pulses whose separations are integer multiples of a very slowly varying period. They are interpreted as pulsars most of whose pulses are nulled. It was easy to show (Katz (2022) used five bursts of FRB 20121102 observed within 93 s by Gajjar et al. (2018)) that RRAT-like periodicity was inconsistent with the bursts of at least one repeating FRB.
The fact that most distributions of intervals between bursts are bimodal may cloud this picture. If the shorter intervals, typically tens of ms, represent substructure within bursts separated by (typically) a few minutes, then it is not possible to define the stable near-exact period required by RRAT-like periodicity.
II.2 Periodic Modulation of Activity
In a magnetar model of FRB the bursts themselves may be stochastic. However, they cannot emit isotropically because a neutron star magnetic field cannot be isotropic. An observation has a detection threshold, and the probability that a burst exceeds that threshold must depend on the orientation of the field with respect to the observer. That probability varies with the rotation period (unless the field is azimuthally symmetric about the rotation axis). This model predicts that the rate of detected bursts varies with the neutron star rotation period. In this note we search for, and set bounds on, such a periodic variation.
III Parameters and Phase Drift
Zhang et al. (2025) observed FRB 20240114A over eight months, which sets a lower bound on its age. No repeating FRB has ever been reported to turn off or even to show a secular decrease in its activity, although their activity varies greatly. This suggests that ages of observed repeating FRB are $\gtrsim 1\,$y, possibly $\gg 1\,$y. FRB 20121102 has been observed to be active for >10>10\,y. Here we assume FRB 20240114A had an age of 1 y when observed on March 12, 2024; numerical results are readily scaled.
A magnetar with a dipole moment component perpendicular to its rotational axis $\mu=\mu_{33}10^{33}\,$Gauss-cm3 (corresponding to an equatorial field of about $10^{15}\,$Gauss), born spinning comparatively rapidly, has an angular rotation frequency
$\omega=\sqrt{{3\over 2}{Ic^{3}\over\mu^{2}A}}\approx 36\mu_{33}^{-1}A_{y}^{-1/2}\text{s}^{-1}$ (1)
[…]
IV Results
[…]
The peridogram of events occuring at times $\{t_{i}\}$ is defined
$P(\omega)=\sqrt{\left(\sum_{i}\cos{(\omega t_{i})}\right)^{2}+\left(\sum_{i}\sin{(\omega t_{i})}\right)^{2}}.$ (2)
It would be possible to weight the terms to account for the strength (flux or fluence) of each burst, but then the small number of very strong bursts would increase the stochastic noise.
Results for frequencies (in cycles/s, not radians) up to 10 Hz are shown in Fig. 1. Extending the search to frequencies up to 100 Hz found no $P(\omega)$ larger than the largest found below 10 Hz (at a frequency of 2.19 Hz). The slight rise of the periodogram to the lowest frequencies indicates a modest variation of the rate of bursts through the 156280 s (4.34 hours) of observation.
Figure 1: Periodogram of 3196 bursts of FRB 20240114A in a 15628 s observation March 12, 2024 (MJD 60381), data from Zhang et al. (2025). Extension of the calculation to frequencies of 100 Hz (implying a neutron star with $\mu_{33}\ll 1$) produced no higher peaks than shown here.
V Frequency Derivative
For a rotor spinning down as a result of its dipole radiation
${\dot{\omega}}=-{\omega\over 2A}\approx-5.7\times 10^{-7}\mu_{33}^{-1}\ A_{y}^{-3/2}\text{s}^{-2}.$ (3)
A large $|\dot{\omega}|$ blurs the signature of periodicity. For an observation of duration $T$ the phase drift of an oscillation from the midpoint of the observation (to which it is fitted) to its ends is
$\Delta\phi={1\over 8}{\dot{\omega}}T^{2}\approx 17\mu_{33}^{-1}A_{y}^{-3/2}\ \text{radian},$ (4)
where for the observation on March 12, 2024 T=15628T=15628\,s. If $A_{y}>10$, $\Delta\phi<0.55\mu_{33}^{-1}\,$radian and spindown during the observation may have little effect on the detectability of periodicity, while if $A_{y}<5$, $\Delta\phi>1.6\mu_{33}^{-1}\,$radian and detectability is siginificantly reduced. Of course, both $A$ and $\mu_{33}$ are unknown a priori. The duration over which the source has been observed sets a lower bound to $A$, as to statistical arguments based on the improbability of discovering a repeating FRB shortly after its birth and arguments based on the radio opacity of very young SNR.
A modified periodogram allows for steady frequency drift:
[…]
where $t_{m}$ is the midpoint of the data range. We calculate this with ${\dot{\omega}}\in[-1\times 10^{-6},+1\times 10^{-6}]\,$s-2 in steps of $10^{-8}$ s-2. Again, no significant peaks were found for any $(\omega,{\dot{\omega}})$ with $2\pi/T\leq\omega\leq 200\pi/$s and the preceding range of $\dot{\omega}$. Yet greater $\omega$ would be inconsistent with the magnetic moment implied by the magnetar hypothesis and an age >1>1\,y.
VI Sensitivity
In order to determine the sensitivity of the periodogram to periodic variation of the rate of FRB we introduce such variation artificially. Bursts are removed from the data with a probability $M\cos{(\omega_{m}t)}$, where $M$ is the modulation amplitude and $\omega_{m}$ the (angular) modulation frequency. We choose a frequency (in cycles) of 1 Hz; the weak dependence of the periodogram on frequency shows that the sensitivity depends only weakly on frequency.
When there are a large number of frequencies considered the clearest way of searching for periodicity is with a bar graph of the distribution of amplitudes (Nowak & Katz (2026) did this in a search for periodicity in the X-ray emission of Sco X-1 in which there were $2^{29}\approx 5.37\times 10^{8}$ frequencies and amplitudes, that could not be readily and legibly plotted). The results are shown in Fig. 2 for several values of $M$.
Figure 2: Distribution of amplitudes of the periodogram of the March 12, 2024 observation (Zhang et al., 2025) of FRB2024014A, with stochastic modulation of various amplitudes $M$ imposed. A 15% modulation of the event rate is readily detected but a 10% modulation cannot be.
VII Discussion
Eq. 3 shows that large negative $\dot{\omega}$ only occurs for neutron stars that are both young and comparatively weakly magnetized; in other words, not magnetars. Quantitatively, our failure to find periodicity (for any possible $\omega$) with ${\dot{\omega}}\in[10^{-6},10^{-6}]$ s-2 indicates $\mu_{33}\leq 0.57A_{y}^{-3/2}$. If A>1A>1\,y, the implied $\mu_{33}$ would be below the range considered to characterize magnetars. Larger $|\dot{\omega}|$ outside the range calculated would require $\omega$ so large as to be inconsistent with Eq. 1 and the assumptions of magnetar fields and A>1A>1\,y.
The possible presence of phase drift is an obstacle to coherent combination of data from observing sessions separated by days. Even if $\dot{\omega}$ were allowed for, over such longer intervals higher derivatives of $\omega$ become significant; measurements of radio pulsars show that these higher derivatives may vary unpredictably.
VIII Conclusion
The magnetar model of FRB makes no quantitative prediction of the magnitude of modulation of the event rate at the magnetar’s rotation frequency. We therefore cannot say that our upper bound of 15% modulation in FRB 20240114A disproves this model. The bound is much lower than the modulation of radiated intensities of known magnetic neutron stars (radio pulsars, accreting neutron star X-ray sources, SGR and X-ray emitting magnetars), that are generally ${\cal O}(100\%)$, but without a quantitative theory of magnetar FRB any inference must be uncertain.
Magnetar models are consistent with tight upper limits on rotational modulation of the rate of detected FRB if either
  1. 1.
    The radiation pattern is azimuthally symmetric about the rotation axis;
  2. 2.
    or, the rotation axis is aligned with the magnetic axis and the line of sight, and emission is collimated along that axis.
The former possibility might require the magnetic field to be nearly azimuthally symmetric ($m=0$ components of all significant multipole orders dominating), perhaps implausible. The latter possibility has the advantage that curvature radiation is aligned with the magnetic axis and may propagate along field lines. Very active repeating FRB may be those with this favorable orientation, so that most of their activity is observable, while apparently non-repeating FRB may be those with mis-aligned axes, so that we only observe their bursts when their magnetic axes are, fortuitously but rarely, aligned with our line of sight.
Eqs. 1 and 3 are inapplicable if the magnetar is accreting. Known magnetars (SGR) are all single objects, but there is no compelling reason why some magnetars could not be born in and remain in mass-transfer binaries; this is why our search included ${\dot{\omega}}>0$. Known mass-transfer binaries containing neutron stars radiate a thermal X-ray spectrum, with no evidence of coherent radio emission (and usually no evidence of particle acceleration at all, although microquasars like Sco X-1 are exceptions). However, all these known accreting neutron stars have magnetic fields below the magnetar range. Accreting magnetars as a source of FRB appear not to have been considered.
Acknowldgements
I thank T. Piran for discussions and the Hebrew University for hospitality during the completion of this work.